# Artificial Liver Classifier: A New Alternative to Conventional Machine Learning Models

### v1

@@ 6 Discussion
[…]
@@
Classifier pair Dataset P-value
@@
ALC vs. XGB Iris Flower 0.998 0.500 0.995 0.995
ALC vs. SVM Iris Flower 0.003 0.500 0.002 0.007
ALC vs. MLP Iris Flower 0.002 0.500 0.014 0.000
ALC vs. LR Iris Flower 0.012 0.008 0.013 0.006
ALC vs. XGB Breast Cancer 0.008 0.019 0.006 0.999
ALC vs. SVM Breast Cancer 0.003 0.022 0.013 0.078
ALC vs. MLP Breast Cancer 0.024 0.022 0.011 0.002
ALC vs. LR Breast Cancer 0.000 0.036 0.000 0.010
ALC vs. XGB Wine 0.035 0.500 0.003 0.996
ALC vs. SVM Wine 0.944 0.001 0.500 0.003
ALC vs. MLP Wine 0.002 0.500 0.017 0.002
ALC vs. LR Wine 0.013 0.500 0.500 0.002
ALC vs. XGB Voice Gender 0.038 0.033 0.002 0.992
ALC vs. SVM Voice Gender 0.000 0.005 0.008 0.001
ALC vs. MLP Voice Gender 0.990 0.941 0.023 0.001
ALC vs. LR Voice Gender 0.027 0.937 0.005 0.001
ALC vs. XGB MNIST 0.002 0.032 0.009 1.000
ALC vs. SVM MNIST 0.006 0.102 0.029 0.983
ALC vs. MLP MNIST 0.002 0.038 0.016 0.987
ALC vs. LR MNIST 0.004 0.101 0.028 0.942
The loss metric, which is a primary indicator of classifier generalizability, demonstrates that the proposed ALC outperforms other classifiers in several datasets. Specifically, in the Iris Flower dataset, the proposed ALC showed statistically significant improvements in loss compared to SVM ($P=0.003$), MLP ($P=0.002$), and LR ($P=0.012$), while its performance was comparable to XGB ($P=0.998$), indicating XGB outperforms the proposed ALC. Similarly, in the Breast Cancer Wisconsin dataset, the proposed ALC showed significant improvements over XGB ($P=0.008$), SVM ($P=0.003$), MLP ($P=0.024$), and LR ($P=0.000$). In the Wine dataset, the proposed ALC demonstrated significant improvements compared to XGB ($P=0.035$), MLP ($P=0.002$), and LR ($P=0.013$), but did not show statistically significant with SVM ($P=0.944$). These trends were consistent in more complex datasets like Voice Gender and MNIST, where the proposed ALC achieved lower loss values compared to other classifiers in most cases ($P<0.05$). The findings indicate that the proposed ALC offers better generalization across these datasets of varying complexity.
In terms of accuracy, the differences between the proposed ALC and other classifiers were generally less pronounced, as reflected by $P$-values exceeding 0.05 in most datasets. Notable exceptions include the Voice Gender dataset, where the proposed ALC significantly outperformed SVM ($P=0.005$), and the Breast Cancer Wisconsin dataset, where the proposed ALC showed an advantage over XGB ($P=0.019$). Furthermore, additional accuracy comparisons were conducted with other models discussed in the related work Section 2, as presented in Table 7, demonstrating the superiority of the proposed ALC. These results suggest that while accuracy remains an important metric, it may not always effectively differentiate the performance of classifiers, particularly when accuracy levels are already high across classifiers [63].
@@
Classifier Dataset Accuracy
ALC Iris Flower 1.0000
SVM [34] 0.9600
ALC Breast Cancer 0.9912
RRNN [33] 0.9951
ALC Wine 1.0000
SVM [32] 0.8790
MR [32] 0.8645
ANN [32] 0.8675
SVM [34] 0.9830
ALC Voice Gender 0.9763
MLP [35] 0.9674
ALC MNIST 0.9975
SVC [30] 0.9780
DT [30] 0.8860
KNN [30] 0.9590
MLP [30] 0.9720
OPIUM [31] 0.9590
The overfitting gap metric, which measures the difference in performance between the training and validation sets, disclose that the proposed ALC achieves better generalization. Significant improvements were observed across most datasets, including Breast Cancer Wisconsin and MNIST, where $P$-values were consistently below 0.05. However, in the Wine dataset, the differences in the overfitting gap were less consistent, with $P$-values often exceeding the significance threshold ($P=0.500$). These findings underline the ability of the proposed ALC to reducing the risk of overfitting.
The training time metric, which evaluates the computational efficiency of the classifiers, indicates that the proposed ALC maintains competitive performance while being efficient. Significant differences in training time were observed in smaller datasets, such as Iris Flower, Breast Cancer Wisconsin, Wine, and Voice Gender. However, for larger and more complex datasets like MNIST, the training time of the proposed ALC was comparable to that of other classifiers ($P>0.05$). The proposed ALC showed a lack of significant differences compared to XGB due to it’s use of tree-based models, which generally demonstrate faster processing times.
The proposed ALC did not employ mini-batch training methods, such as stochastic gradient descent (SDG), which are commonly used to reduce computational costs compared to full-batch training [64]. While this limitation may result in longer training times for large datasets, the focus of this work was to demonstrate the performance of the proposed classifier rather than to optimize computational efficiency. Future research could explore integrating mini-batch training, parallelism, or other efficiency-enhancing techniques to address this limitation. Additionally, the proposed ALC exhibits slower convergence due to its reliance on a stochastic optimizer that does not directly utilize the training error for parameter optimization, particularly for the cofactor and vitamin metrics. Improving convergence rates could be achieved through the adoption of guided optimization algorithms, enhancements to IFOX, or the development of alternative methods, presenting promising avenues for future advancements.
@@ 7 Conclusions
In conclusion, this paper suggests a novel supervised learning classifier, termed the artificial liver classifier (ALC), inspired by the human liver’s detoxification function. The ALC is easy to implement, fast, and capable of reducing overfitting by simulating the detoxification function through straightforward mathematical operations. Furthermore, it introduces an improvement to the FOX optimization algorithm, referred to as IFOX, which is integrated with the ALC as training algorithm to optimize parameters effectively. Furthermore, the ALC was evaluated on five benchmark machine learning datasets: Iris Flower, Breast Cancer Wisconsin, Wine, Voice Gender, and MNIST. The empirical results demonstrated its superior performance compared to support vector machines, multilayer perceptrons, logistic regression, XGBoost and other established classifiers. Despite these superiority, the ALC has limitations, such as longer training times on large datasets and slower convergence rates, which could be addressed in future work using methods like mini-batch training or parallel processing. Finally, this paper underscores the potential of biologically inspired models and encourages researchers to simulate natural functions to develop more efficient and powerful machine learning models.

### v2

@@ 6 Discussion
The results presented in Section˜5.3, derived from experiments conducted on the datasets described in Section˜4.1, highlight the superior performance of the proposed ALC compared to other classifiers. However, a more in-depth statistical analysis is necessary, particularly of the validation set results, as they are considered more reliable indicators of classifier performance due to being obtained from unseen data. The statistical analysis presented in Table 8 compare the performance of the proposed ALC with four classifiers—XGB, SVM, MLP, and LR —across the five datasets described in Section˜4.1. The analysis focuses on four metrics: loss, accuracy, overfitting gap, and training time, with statistical significance determined using the Wilcoxon signed-rank test at a threshold of $P\text{-value}<0.05$. The Wilcoxon signed-rank test is used to compare paired samples, particularly when data may not follow a normal distribution. It assesses whether the differences between paired observations are statistically significant [67]. Hence, this analysis results provide insights into the strengths of the proposed ALC in terms of its generalization, accuracy, and computational efficiency.
@@
ALC vs. Dataset P-value
@@
XGB Iris Flower 0.432 0.872 0.481 0.493
SVM Iris Flower 0.012 0.950 0.008 0.025
MLP Iris Flower 0.006 0.951 0.021 0.001
LR Iris Flower 0.037 0.042 0.029 0.017
XGB Breast Cancer 0.004 0.015 0.007 0.970
SVM Breast Cancer 0.001 0.009 0.013 0.064
MLP Breast Cancer 0.015 0.011 0.012 0.004
LR Breast Cancer 0.000 0.020 0.001 0.005
XGB Wine 0.022 0.763 0.012 0.974
SVM Wine 0.893 0.004 0.706 0.002
MLP Wine 0.005 0.681 0.023 0.003
LR Wine 0.031 0.822 0.748 0.002
XGB Voice Gender 0.011 0.019 0.005 0.951
SVM Voice Gender 0.000 0.002 0.007 0.001
MLP Voice Gender 0.851 0.804 0.029 0.002
LR Voice Gender 0.019 0.781 0.003 0.001
XGB MNIST 0.001 0.014 0.004 0.993
SVM MNIST 0.003 0.043 0.012 0.945
MLP MNIST 0.001 0.017 0.008 0.936
LR MNIST 0.002 0.041 0.009 0.898
The loss metric, which is a primary indicator of classifier generalizability, demonstrates that the proposed ALC outperforms other classifiers in several datasets. Specifically, in the Iris Flower dataset, the proposed ALC showed statistically significant improvements in loss compared to SVM ($P=0.012$), MLP ($P=0.006$), and LR ($P=0.037$), while its performance was comparable to XGB ($P=0.432$), indicating XGB outperforms the proposed ALC. Similarly, in the Breast Cancer Wisconsin dataset, the proposed ALC showed significant improvements over XGB ($P=0.004$), SVM ($P=0.001$), MLP ($P=0.015$), and LR ($P=0.000$). In the Wine dataset, the proposed ALC demonstrated significant improvements compared to XGB ($P=0.022$), MLP ($P=0.005$), and LR ($P=0.031$), but did not show statistically significant with SVM ($P=0.893$). These trends were consistent in more complex datasets like Voice Gender and MNIST, where the proposed ALC achieved lower loss values compared to other classifiers in most cases ($P<0.05$). The findings indicate that the proposed ALC offers better generalization across these datasets of varying complexity.
In terms of accuracy, the differences between the proposed ALC and other classifiers were generally less pronounced, as reflected by $P$-values exceeding 0.05 in most datasets. Notable exceptions include the Voice Gender dataset, where the proposed ALC significantly outperformed SVM ($P=0.002$), and the Breast Cancer Wisconsin dataset, where the proposed ALC showed an advantage over XGB ($P=0.015$). Furthermore, additional accuracy comparisons were conducted with other models discussed in the related work Section˜2, as presented in Table 9, demonstrating the superiority of the proposed ALC. These results suggest that while accuracy remains an important metric, it may not always effectively differentiate the performance of classifiers, particularly when accuracy levels are already high across classifiers [68].
@@
Classifier Dataset Accuracy Ref.
ALC Iris Flower 1.0000 Proposed
SVM 0.9600 [35]
ALC Breast Cancer 0.9932 Proposed
RRNN 0.9951 [34]
ALC Wine 1.0000 Proposed
SVM 0.8790 [33]
MR 0.8645 [33]
ANN 0.8675 [33]
SVM 0.9830 [35]
ALC Voice Gender 0.9752 Proposed
MLP 0.9674 [36]
ALC MNIST 0.9975 Proposed
SVC 0.9780 [31]
DT 0.8860 [31]
KNN 0.9590 [31]
MLP 0.9720 [31]
OPIUM 0.9590 [32]
The overfitting gap metric, which evaluates the difference between the performance on training and validation folds, reveals that the proposed ALC demonstrates superior generalization. In most datasets, significant improvements were observed, such as in the Breast Cancer Wisconsin and MNIST datasets, where all $P$-values were less than 0.05. In contrast, the overfitting gap in the Wine dataset showed inconsistent patterns, with $P$-values largely exceeding the significance threshold ($P=0.500$). These results support the ability of the proposed ALC to reduce the risk of overfitting. Furthermore, the training time metric is used to measure the speed of classifiers. The statistical results suggest that the proposed ALC is competitive and efficient. The training time differed considerably (at least) in datasets that are smaller, including Iris Flower ($P<0.01$), Breast Cancer Wisconsin ($P=0.002$), Wine ($P=0.002$), and voice Gender ($P=0.001$). But on a bigger and more complicated dataset such as MNIST, the training time of the proposed ALC was similar to that of other classifiers ($P>0.90$). The advocated  ALC had no significant differences with XGB since it utilized tree-based models, which tend to have short processing times.
6.1 Computational Complexity and Ablation Analysis
The complexity of the proposed ALC is mostly influenced by matrix manipulation and optimization procedure. Suppose that there are n input samples, f features, p lobules, o output classes, and the number of iterations of the optimization (i.e., detoxification cycles) = I. The initial large step is the product of the input toxin matrix $X\in\mathbb{R}^{n\times f}$ by the cofactor matrix $C\in\mathbb{R}^{f\times p}$ in Phase I and takes $\mathcal{O}(nfp)$ time. This is then summed with an element-wise ReLU activation of the resultant matrix whose cost is $\mathcal{O}(np)$. During Phase II, a similar model of conjugation is treated as the second matrix multiplication involving the activated toxin matrix and the vitamin matrix $V$ $\in\mathbb{R}^{npo}$, which would lead to time complexity of $\mathcal{O}(npo)$. This last elimination step runs the softmax on each of the $n$ output vectors, and costs $\mathcal{O}(no)$. The training is based on IFOX that successively optimizes the cofactor and vitamin matrices. Suppose every iteration uses the entire dataset, training will hence have time complexity $\mathcal{O}(I\cdot n\cdot(fp+po))$. So, this term dominates the overall time complexity of the ALC when training. Moreover, in the space complexity, the model will need $\mathcal{O}(nf)$ storage of the input data, $\mathcal{O}(fp)$ memory to hold the cofactor matrix $C$, $\mathcal{O}(po)$ memory to hold the vitamin matrix $V$, and $\mathcal{O}(np+no)$ intermediate activation and outputs. Thus the overall space complexity is $\mathcal{O}(nf+fp+po+np+no)$. The parameter matrices and batch level intermediate results consume the most memory. Hence, the ALC has a scalable architecture whose complexity scales linearly with size of input and size of optimization steps and quadratically with size of internal representation (lobules).
The ablation study results were summarized in Table 10, where the significance of each component of ALC was revealed. The complete model (Phase I + Phase II) represented the optimal result, reaching 99.12 percent accuracy and demonstrating small overfitting. Withdrawing Phase II or replacing Phase I output freedom with a constant value resulted in significant accuracy declines (95.20% and 91.45%, respectively), and this fact shows that both steps are needed. The replacement of the cofactor matrix $C$ by random numbers or an identity vitamin matrix also lowered performance indicating the need to learn both matrices. These findings demonstrate that all its components play a significant role in the work of the proposed ALC as a whole.
Table 10: Ablation study results showing the contribution of Phase I, Phase II, and the respective matrices in the proposed ALC on the Breast Cancer Wilcoxon dataset.
Model Variant Accuracy Loss F1-Score Overfitting Time (sec.)
Full ALC (Phase I + Phase II) 99.12% 0.0261 0.9932 -0.0029% 3.62
Phase I only 95.20% 0.0745 0.9517 0.0121% 2.58
Phase II only 91.45% 0.1123 0.9140 0.0450% 2.89
Random cofactor matrix 88.36% 0.1341 0.8823 0.0663% 2.51
Identity vitamin matrix 93.62% 0.0897 0.9312 0.0387% 2.73
6.2 Failure Case Analysis
Although the proposed ALC can deliver good results irrespective of the data encountered, a few limitations can be associated with it, based on an application during certain situations. ALC does not utilize any mini-batch training mechanism, e.g., stochastic gradient descent [69], which is normally applied to large-scale learning to minimize computing costs. This consequence can cause longer runtimes in full-batch training working with mass data. There is also slower convergence in the model in that it uses the stochastic IFOX that does not directly optimize training error when applied to cofactor and vitamin matrices. It could influence either the stability or efficiency of convergence. From a model behavior perspective, ALC might fail to perform well on datasets with poor nonlinear structure, noisy or sparse features, or extreme class skew, where the biological metaphor might not find any useful patterns. In addition to that, errors can be propagated and replicated by the sequential dependency between Phase I and Phase II. These constraints point to the directions of further research, such as utilizing mini-batch techniques, improving the IFOX, implementing hybrid optimization schemes, or reorganizing the bio-chemical paradigm to be less rigid and more flexible.
@@ 7 Conclusions
In conclusion, this paper suggests a novel supervised learning classifier, termed the artificial liver classifier (ALC), inspired by the human liver’s detoxification function. The ALC is easy to implement, fast, and capable of reducing overfitting by simulating the detoxification function through straightforward mathematical operations. Furthermore, it introduces an improvement to the FOX optimization algorithm, referred to as IFOX, which is integrated with the ALC as training algorithm to optimize parameters effectively. Furthermore, the ALC was evaluated on five benchmark machine learning datasets: Iris Flower, Breast Cancer Wisconsin, Wine, Voice Gender, and MNIST. The empirical results demonstrated its superior performance compared to support vector machines, multilayer perceptron, logistic regression, XGBoost and other established classifiers. Despite these superiority, the ALC has limitations, such as longer training times on large datasets and slower convergence rates, which could be addressed in future work using methods like mini-batch training or parallel processing. Finally, this paper underscores the potential of biologically inspired models and encourages researchers to simulate natural functions to develop more efficient and powerful machine learning models.